OpenStudy (lukecrayonz):

Why is the integral of 17e^t, 17e^t?

OpenStudy (ikram002p):

its 17e^t... do integration by parts and ul gonna figer out. can u do that ??

OpenStudy (ikram002p):

ok skip "by parts method" u know that the integration is the opposite of derivative. so Asuume u have U =e^t du=17e^t intg du = intg 17 e^t =17(ing e^t)//// by difintion.... intg e^t= U+c right ?? =17(U+c) =17U+c1 =17e^t+c1 ..